One number is bigger than another by 406. If the bigger number is divided by the smaller one, you’ll get the quotient 3 and remainder 66. Find the numbers.
Answer by ikleyn(52776)

.
The condition means


    x + 406 = 3x + 66.


Solve for x, which stands for the smaller number


    406 - 66 = 3x - 2x

    340 = 2x   ====================>  x = 340/2 = 170.


ANSWER.  The smaller number is 170;  the larger one is  170 + 406 = 576.

You can check my answer on your own.